Give the steps you will use to separate the variable and then solve the equation:
20p/3 = 40

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

First we have to multiply both sides of the equation by 3,
Then, we get,
= (20p/3) × 3 = 40 × 3
= 20p = 120
Now,
We have to divide both sides of the equation by 20,
Then, we get,
= 20p/20 = 120/20
= p = 6

Thanks